calculate the area of a circular disc 0.86m in diameter. (π=3.142)
1
Expert's answer
2020-10-15T18:47:22-0400

The area "A" of a circular disk with diameter "d" is


"A=\dfrac{\pi\cdot d^2 }{4}"

The area of a circular disc 0.86m in diameter is


"A=\dfrac{\pi\cdot(0.86m)^2}{4}=0.1849\pi m^2\approx0.5809558m^2"

The area of a circular disc 0.86m in diameter is "0.5809558m^2."